# A Simple "Motivation" Can Enhance Reinforcement Finetuning of Large Reasoning Models

### v2

@@ Abstract
Figure 1: Validation Accuracy of MeRF and RLVR baseline on K&K Logic Puzzles. Simply “telling LLMs rules of the game” with in-context motivation during training, MeRF significantly outperforms the RLVR baseline with better efficiency, demonstrating the effectiveness of leveraging in-context motivation for more efficient RL training.
@@ 1 Introduction
Large Language Models (LLMs) have demonstrated remarkable capabilities across a wide range of natural language understanding and generation tasks, such as instruction following (Ouyang et al., 2022; Zhou et al., 2023; Taori et al., 2023), code generation (Chen et al., 2021; Nijkamp et al., 2022; Zhuo et al., 2024), and medical diagnosis (Singhal et al., 2025; Zhang et al., 2023; Wang et al., 2023). To further improve the reasoning capabilities of LLMs, Reinforcement Learning with Verifiable Rewards (RLVR) (Lambert et al., 2024; Team et al., 2025) has emerged as a promising alternative to conventional supervised fine-tuning approaches (Radford et al., 2018; Brown et al., 2020), as demonstrated by DeepSeek-R1 (Guo et al., 2025) and OpenAI-o1 (Jaech et al., 2024). RLVR treats reasoning as a sequential decision-making process and optimizes models using objective reward signals that can be automatically verified with explicit rules, such as matching mathematical answers to ground truth (Lambert et al., 2024; He et al., 2024), or checking code correctness through unit tests (Le et al., 2022; Liu et al., 2023). By optimizing models toward meeting the clearly defined success criteria, RLVR enables LLMs to iteratively refine their reasoning capabilities, achieving significant performance improvements on complex tasks.
Following and extending the success of the conventional RL paradigm in gaming, Go, and robotics to the field of language models, RLVR relies on external reward signals to guide the training process of LLMs (Zeng et al., 2025; Yu et al., 2025; Hu et al., 2025; Chen et al., 2025). However, in the same way as conventional reinforcement learning, RLVR is not very efficient, as it works in a trial-and-error manner, where the model can suffer from the sparse reward space and needs to learn the patterns of the task and reward function by repeatedly collecting and comparing its outputs and corresponding rewards. It requires a larger amount of training data and computational resources for LLMs to learn from and generate enough responses to reach the expected behavior and then get positive feedback. As we continue expecting LLMs to solve more and more complex tasks, the reward function can be more sophisticated and the expected behavior can be harder to reach, making it even more important to improve the efficiency of the current RLVR paradigm.
@@ 2 Method
Figure 2: (Left) Illustration of the RLVR pipeline and the in-context motivation introduced by MeRF. Compared to the indirect way (reward samples generated and through parameter updates) to learn the reward patterns, MeRF enables the model to be aware of the overall reward space by in-context motivation. (Right) We validate models in two settings: w/ motivation and w/o motivation in the prompt. Different from the base model, the RLVR model achieves a slightly better performance in validation w/ motivation than w/o motivation after the RLVR training, even while the motivation is not involved in the training process, indicating a connection between the in-context motivation validation and the RLVR training guided by the reward function (as the motivation describes).
@@ 2.1 Motivation-enhanced Reinforcement Finetuning
In Figure 2 (Right), we conduct an experiment, comparing the performance of the Base model and RLVR model with and without motivation in the inference time, where we find that the Base Model does not benefit from the motivation in the inference time, while the RLVR model, trained with the reward function described in the motivation, achieves an improvement when validated with in-context motivation. Moreover, when we involve the motivation during the training process (MeRF), we find a significant performance improvement over the RLVR baseline, demonstrating the effectiveness of MeRF in improving the efficiency and effectiveness of reinforcement finetuning of LLMs.
@@ 2.2 Motivation with Verifiable Reward
<|im_start|>user {input the puzzle} <|im_end|>
<|im_start|>assistant <think>
By injecting the motivation into the training process, we enable the model to be aware of the motivation of the task, which describes the reward function of the RLVR pipeline. The motivation provides a clear specification of what is expected and how to do it correctly, aligning the generation with the transcendent optimization objective. This approach leverages the in-context learning ability of LLMs to improve their reasoning capabilities in a more efficient and effective manner.
@@ 3.2 Main Results
To demonstrate the effectiveness of MeRF in the reinforcement finetuning process, we compare the performance of MeRF with the RLVR baseline on the K&K Logic Puzzle. The results in Figure 1 and Figure 3 show that MeRF consistently achieves a significant improvement over the RLVR baseline in the validation accuracy during the training process, across different model sizes and model families, revealing the remarkable effectiveness of the motivation-enhanced reinforcement finetuning, with simply injecting the in-context motivation.
@@
Figure 3: Validation Accuracy of MeRF and RLVR baseline on two DeepSeek-R1-Distill models of different sizes and from different model families on K&K. MeRF consistently outperforms the RLVR baseline, demonstrating the effectiveness and generalization of MeRF.
Table 2: Comparison between RLVR baseline and MeRF on math reasoning datasets. We report pass@$k$ for $k\!\in\!\{1,2,4,8\}$ on AIME24&25, AMC23, and MATH500 datasets. Best results are in bold; Average block highlights overall performance and shows absolute gains over Baseline in parentheses. MeRF consistently outperforms the RLVR baseline across these datasets and metrics.
Method AIME24 AIME25 AMC23
pass@1 pass@2 pass@4 pass@8 pass@1 pass@2 pass@4 pass@8 pass@1 pass@2 pass@4 pass@8
RLVR (Baseline) 16.7 20.0 20.0 26.7 6.7 16.7 16.7 20.0 47.5 57.5 70.0 72.5
MeRF (Ours) 20.0 20.0 26.7 30.0 6.7 10.0 16.7 26.7 55.0 67.5 72.5 77.5
@@ 4 Analysis on the mechanism behind MeRF
We conduct experiments to investigate where the performance improvement mainly comes from. As shown in Figure 2 (Right), we observe that for both RLVR and MeRF model, motivation validation leads to slightly better performance (4% and 2%) than non-motivation validation. However, compared to the performance improvement of MeRF over RLVR in both non-motivation validation and motivation validation (27% and 25%), the performance improvement of MeRF is mainly from the motivation-enhanced reinforcement finetuning process, demonstrating the effectiveness of the in-context motivation in the training process. The results in Figure 8 also show that the performance of both models does not benefit much from the in-context motivation validation, indicating that the performance improvement of MeRF is not from the in-context inference.
@@
To further understand how the in-context motivation helps the reinforcement finetuning process, we analyze the pass@$k$ performance and entropy of the models during the training process. The results in Figure 5 show that MeRF consistently outperforms the RLVR baseline in all pass@$k$ metrics, demonstrating the effectiveness of MeRF in improving the reasoning capabilities of LLMs. More importantly, while RLVR’s performance of pass@4 and pass@8 hardly improves after step 140, MeRF achieves better pass@4 and pass@8 performance at step 140 than the final RLVR Model (at step 280). The results in Figure 7 also show that MeRF outperforms the RLVR baseline consistently in both pass@8 and pass@1 metrics on the MATH500 dataset, while RLVR pass@8 performance hardly improves after step 80. Pass@$k$ metrics serve as an indicator of the model’s ability to explore diverse reasoning paths and reach a correct answer (Zeng et al., 2025; Shao et al., 2024). Better pass@$k$ performance and continuous pass@$k$ improvement during the training process suggest the model is more likely to reach a positive reward for optimization and performance improvement. As 8 is the rollout group size of GRPO in our experiments, the growing improvement of pass@8 performance of MeRF over the RLVR baseline indicates that training process progressively amplifies the initial pass@8 improvement with better exploration ability, initially benefiting from the in-context motivation, which also explains why the motivation validation only leads to slightly better performance than non-motivation validation but significant performance improvement of MeRF over RLVR baseline in Figure 2 (Right). The results in Figure 4 show that MeRF maintains a higher entropy than the RLVR baseline during the training process, indicating that MeRF encourages more exploration by the in-context motivation, avoiding the entropy collapse issue of RL for LLMs (Cui et al., 2025), which contributes to its improved performance. The results in Figure 6 also show that MeRF outperforms the RLVR baseline with better accuracy and pass@4 performance on the CountDown dataset, while maintaining a higher entropy during the training process. The results suggest that MeRF benefits from the exploration ability encouraged by the in-context motivation, leading to better performance in the reinforcement finetuning process.
@@
Figure 8: Validation accuray w/ motivation and w/o motivation in the prompt. Both models achieve comparable performance between two validation settings, indicating the negligible impact of the training and validation gap.
@@
The results demonstrate that the motivation with ground-truth reward function achieves the best performance, and the suboptimal motivation performs better than the RLVR baseline, with an additional correctness score description included in the motivation. Adverse motivation provides the full description of the reward function same to ground-truth motivation, while all the score is assigned to the opposite, which misleads the model to provide the wrong answer. The performance drop of the adverse motivation is mainly caused by the contradiction between the in-context motivation and the reward function of the optimization process. However, after several rounds of unstable learning dynamics, the model adapted to discount the motivation signal, either treating the score as uninformative, or understanding the deliberately adverse motivation and opposite scores, while the full description of the reward function is still beneficial for the model to outperform the RLVR baseline and Motivation (Suboptimal). The results suggest that MeRF benefits from a better consistency between the motivation and the underlying reward function, resulting in better performance, while the model is capable of adapting to the adverse motivation in the process of reinforcement finetuning.
@@ 5 Conclusion
Reproducibility Statement
We have taken several steps to ensure the reproducibility of our work. The main paper provides a complete description of the proposed method and experimental setup. The appendix includes additional details on the training procedures, and hyperparameter settings. We will release the code, and detailed instructions for reproducing our experiments upon acceptance.
@@ A.1 Use of LLMs
A.2 Related Work
Reinforcement Learning for LLMs. Reinforcement learning (RL) has become a powerful paradigm for fine-tuning LLMs, first demonstrated by reinforcement learning from human feedback (RLHF) to align models with human-preferred responses (Ouyang et al., 2022; Rafailov et al., 2023). More recently, inspired by the success of DeepSeek-R1 (Guo et al., 2025), Reinforcement Learning with Verified Reward (RLVR) (Lambert et al., 2024) has been proposed to directly enhance reasoning performance by rewarding verifiable success criteria instead of subjective human judgments, which is more similar to the traditional RL paradigm. RLVR has shown promising results in improving the performance of LLMs on various reasoning tasks, including logical reasoning (Xie et al., 2025), coding (Le et al., 2022; Liu et al., 2023), and math problems (Zeng et al., 2025), with simple, designed reward functions. However, these approaches demonstrate the potential of the conventional RL paradigm for LLMs, which may not fully leverage the in-context learning abilities of LLMs, the key to the success of LLMs in previous works (Wei et al., 2022; Brown et al., 2020). In this work, we propose a novel method MeRF, leveraging the in-context abilities of LLMs to enhance reinforcement finetuning for reasoning.
A.3 Preliminary

### v3

@@ Abstract
Figure 1: Validation Accuracy of MeRF and RLVR baseline on K&K Logic Puzzles in the training process. By simply “telling LLMs rules of the game” with in-context motivation during RL training, MeRF significantly outperforms the RLVR baseline with faster improvements, demonstrating the effectiveness of leveraging in-context motivation for more efficient RL training of LLMs.
@@ 1 Introduction
Large Language Models (LLMs) have demonstrated remarkable capabilities across a wide range of natural language understanding and generation tasks, such as instruction following (Ouyang et al., 2022; Zhou et al., 2023; Taori et al., 2023; Zhu et al., 2024; Tu et al., 2025b; Zhu et al., 2025), code generation (Chen et al., 2021; Nijkamp et al., 2022; Zhuo et al., 2024), and medical diagnosis (Singhal et al., 2025; Zhang et al., 2023; Wang et al., 2023). To further improve the reasoning capabilities of LLMs, Reinforcement Learning with Verifiable Rewards (RLVR) (Lambert et al., 2024; Team et al., 2025) has emerged as a promising alternative to conventional supervised fine-tuning approaches (Radford et al., 2018; Brown et al., 2020; Zhang et al., 2025a), as demonstrated by DeepSeek-R1 (Guo et al., 2025) and OpenAI-o1 (Jaech et al., 2024). RLVR treats reasoning as a sequential decision-making process and optimizes models using objective reward signals that can be automatically verified with explicit rules, such as matching mathematical answers to ground truth (Lambert et al., 2024; He et al., 2024), or checking code correctness through unit tests (Le et al., 2022; Liu et al., 2023). By optimizing models toward meeting the clearly defined success criteria, RLVR enables LLMs to iteratively refine their reasoning capabilities, achieving significant performance improvements on complex tasks.
Following and extending the success of the conventional RL paradigm in gaming, Go, and robotics to the field of language models, RLVR relies on external reward signals to guide the training process of LLMs (Zeng et al., 2025; Yu et al., 2025; Hu et al., 2025; Chen et al., 2025; Zhang et al., 2025b; Fang et al., 2025). However, in the same way as conventional reinforcement learning, RLVR is not efficient enough, as it works in a trial-and-error manner, where the model can suffer from the sparse reward space and needs to learn the patterns of the task and reward function by repeatedly collecting and comparing its outputs and corresponding rewards. It requires a larger amount of training data and computational resources for LLMs to learn and generate enough responses to reach the expected behavior, and then get positive feedback. As we continue expecting LLMs to solve more and more complex tasks, the reward function can be more sophisticated and the expected behavior can be harder to reach, making it even more important to improve the efficiency of the current RLVR paradigm.
@@
2 Related Work
Reinforcement Learning for LLMs. RL has become a powerful paradigm for fine-tuning LLMs, first demonstrated by reinforcement learning from human feedback (RLHF) to align models with human-preferred responses (Ouyang et al., 2022; Rafailov et al., 2023; Liu et al., 2025). More recently, inspired by the success of DeepSeek-R1 (Guo et al., 2025), Reinforcement Learning with Verified Reward (RLVR) (Lambert et al., 2024) has been proposed to directly enhance reasoning performance by rewarding verifiable success criteria instead of subjective human judgments, which is more similar to the traditional RL paradigm (Jiang et al., 2021; 2022). RLVR has shown promising results in improving the performance of LLMs on various reasoning tasks, including logical reasoning (Xie et al., 2025), coding (Le et al., 2022; Liu et al., 2023), and math problems (Zeng et al., 2025), with simple, designed reward functions. However, these approaches demonstrate the potential of the conventional RL paradigm for LLMs, which may not fully leverage the in-context learning abilities of LLMs, the key to the success of LLMs in previous works (Wei et al., 2022; Brown et al., 2020; Tu et al., 2025a). In this work, we propose a novel method MeRF, leveraging the in-context abilities of LLMs to enhance reinforcement finetuning for reasoning.
@@ 3 Method
Figure 2: (Left) Illustration of the RLVR pipeline and the in-context motivation introduced by MeRF. Compared to the indirect way (reward samples generated and through parameter updates) to learn the reward patterns, MeRF enables the model to be aware of the overall reward space by in-context motivation. (Right) We validate the Base model, RLVR model and MeRF model on the K&K Logic Puzzle dataset in two settings: w/ motivation and w/o motivation in the prompt. Different from the base model, the RLVR model achieves a slightly better performance in validation w/ motivation than w/o motivation after the RLVR training, even while the motivation is not involved in the training process, indicating a connection between the in-context motivation validation and the RLVR training guided by the reward function (as the motivation describes).
@@ 3.1 Motivation-enhanced Reinforcement Finetuning
In Figure 2 (Right), we conduct an experiment to investigate how the in-context motivation affects the performance in the inference process and training process respectively. First, comparing the performance of the Base model with and without motivation in the inference time, Base model does not benefit from the motivation, indicating that the model can’t learn the reward patterns and expected behavior by in-context motivation alone. Next, RLVR model (trained with the reward function described in the motivation), achieves an improvement when validated with in-context motivation. Moreover, when we involve the motivation during the training process (MeRF), there is a significant performance improvement over the RLVR baseline, demonstrating the effectiveness of MeRF in improving the efficiency and effectiveness of reinforcement finetuning of LLMs.
@@ 3.2 Motivation with Verifiable Reward
By injecting the motivation into the training process, we enable the model to be aware of the motivation of the task, which describes the reward function of the RLVR pipeline. The motivation provides a clear specification of what is expected and how to do it correctly, aligning the generation with the transcendent optimization objective. This approach leverages the in-context learning ability of LLMs to improve their reasoning capabilities in a more efficient and effective manner. More details about the complete motivation prompt design for the tasks are provided in Appendix A.3.
@@ 4.1 Experimental Setup
Qwen2.5-7B-Instruct-1M 0.40 0.25 0.11 0.06 0.02 0.17 0.49 0.01 0.19
   +RLVR (Baseline) 0.52 0.44 0.36 0.18 0.12 0.32 0.60 0.16 0.34
[…]
Qwen2.5-14B-Instruct 0.46 0.31 0.20 0.09 0.11 0.23 0.63 0.06 0.27
   +RLVR (Baseline) 0.90 0.82 0.78 0.70 0.55 0.75 0.90 0.42 0.72
[…]
DeepSeek-R1-Distill-Qwen-1.5B 0.08 0.01 0.00 0.00 0.00 0.02 0.30 0.00 0.06
   +RLVR (Baseline) 0.20 0.16 0.16 0.10 0.04 0.13 0.29 0.01 0.14
[…]
DeepSeek-R1-Distill-Llama-8B 0.26 0.22 0.14 0.06 0.05 0.15 0.24 0.11 0.15
   +RLVR (Baseline) 0.92 0.89 0.84 0.79 0.64 0.82 0.90 0.58 0.79
[…]
@@ 4.2 Main Results
To demonstrate the effectiveness of MeRF in the reinforcement finetuning process, we compare the performance of MeRF with the RLVR baseline on the K&K Logic Puzzle. The results in Figure 1 show that MeRF consistently achieves a significant improvement over the RLVR baseline in the validation accuracy during the training process, across different model sizes and model families, revealing the remarkable effectiveness of the motivation-enhanced reinforcement finetuning, with simply injecting the in-context motivation.
@@
Table 2: Comparison of RLVR baseline and MeRF across math reasoning datasets. Each dataset occupies one row, with both methods displayed in vertical blocks (pass@1/2/4/8). The last row summarizes the average performance.
Dataset RLVR (Baseline) MeRF (Ours)
pass@1 pass@2 pass@4 pass@8 pass@1 pass@2 pass@4 pass@8
AIME24 16.7 20.0 20.0 26.7 20.0 20.0 26.7 30.0
AIME25 6.7 16.7 16.7 20.0 6.7 10.0 16.7 26.7
AMC23 47.5 57.5 70.0 72.5 55.0 67.5 72.5 77.5
MATH500 62.6 72.8 77.0 82.6 65.4 74.0 81.8 85.6
@@ 5 Analysis on the mechanism behind MeRF
To answer this question, we conduct experiments to investigate how in-context motivation affects the performance in inference process and training process. As shown in Figure 2 (Right), we observe that for both RLVR and MeRF model, motivation validation leads to slightly better performance (4% and 2%) than non-motivation validation. However, compared to the performance improvement of MeRF over RLVR in both non-motivation validation and motivation validation (27% and 25%), the performance improvement of MeRF is mainly from the motivation-enhanced reinforcement finetuning process, demonstrating the effectiveness of the in-context motivation in the training process. The results in Figure 7 also show that the performance of both models does not benefit much from the in-context motivation validation, indicating that the performance improvement of MeRF is not from the in-context inference.
@@
To further understand how the in-context motivation helps the reinforcement finetuning process, we analyze the pass@$k$ performance and entropy of the models during the training process. The results in Figure 3 show that MeRF consistently outperforms the RLVR baseline in all pass@$k$ metrics, demonstrating the effectiveness of MeRF in improving the reasoning capabilities of LLMs. More importantly, while RLVR’s performance of pass@4 and pass@8 hardly improves after step 140, MeRF achieves better pass@4 and pass@8 performance at step 140 than the final RLVR Model (at step 280). The results in Figure 4 also show that MeRF outperforms the RLVR baseline consistently in both pass@8 and pass@1 metrics on the MATH500 dataset, while RLVR pass@8 performance hardly improves after step 80. Pass@$k$ metrics serve as an indicator of the model’s ability to explore diverse reasoning paths and reach a correct answer (Zeng et al., 2025; Shao et al., 2024). Better pass@$k$ performance and continuous pass@$k$ improvement during the training process suggest the model is more likely to reach a positive reward for optimization and performance improvement. As 8 is the rollout group size of GRPO in our experiments, the growing improvement of pass@8 performance of MeRF over the RLVR baseline indicates that training process progressively amplifies the initial pass@8 improvement with better exploration ability, initially benefiting from the in-context motivation, which also explains why the motivation validation only leads to slightly better performance than non-motivation validation but significant performance improvement of MeRF over RLVR baseline in Figure 2 (Right).
Figure 5: Correct Ratio of generated answers of the training set during the training process on K&K Logic Puzzle dataset. MeRF consistently outperforms the RLVR baseline, demonstrating the better exploration ability encouraged by the in-context motivation for the model to get the best reward during the training process.
As the result shown in Figure 6, MeRF maintains a higher entropy than the RLVR baseline. Specifically, we find that at the beginning of training, MeRF has lower entropy than RLVR, indicating that the structured exploration introduced by motivations helps the model to focus on more promising areas of the output space compared with naive exploration. As training progresses, RLVR entropy decreases more rapidly, indicating that the model is losing exploration ability and somehow collapsing (Cui et al., 2025) to suboptimal solutions (e.g., only get format reward) due to sparse reward signal and blind exploration. In contrast, MeRF maintains a relatively higher entropy throughout training, suggesting that the reward space information provided by motivations encourages the model to explore more effectively for the best reward (correct answer, as shown in Figure 5) rather than collapsing to suboptimal solutions rapidly.
@@
Figure 7: Validation Accuray w/ motivation and w/o motivation in the prompt for MeRF Models. The comparable performance between two validation settings of both MeRF Models, indicating the negligible impact of the training and validation gap in terms of the prompt, which does not necessarily lead to a drop in performance.
@@
The results demonstrate that the motivation with ground-truth reward function achieves the best performance, and the suboptimal motivation performs better than the RLVR baseline, with an additional correctness score description included in the motivation. Adverse motivation provides the full description of the reward function same to ground-truth motivation, while all the score is assigned to the opposite, which misleads the model to provide the wrong answer. The performance drop of the adverse motivation is mainly caused by the contradiction between the in-context motivation and the reward function of the optimization process. However, after several rounds of unstable learning dynamics (more results are provided in the Figure 10), the model adapted to discount the motivation signal, either treating the score as uninformative, or understanding the deliberately adverse motivation and opposite scores, while the full description of the reward function is still beneficial for the model to outperform the RLVR baseline and Motivation (Suboptimal). The results suggest that MeRF benefits from a better consistency between the motivation and the underlying reward function, resulting in better performance, while the model is capable of adapting to the adverse motivation in the process of reinforcement finetuning.
@@ 6 Conclusion
Acknowledgments
This research is supported by the RIE2025 Industry Alignment Fund – Industry Collaboration Projects (IAF-ICP) (Award I2301E0026), administered by A*STAR, as well as supported by Alibaba Group and NTU Singapore through Alibaba-NTU Global e-Sustainability CorpLab (ANGEL).
@@ A.1 Use of LLMs
A.2 Preliminary
@@
A.3 Prompt Design of In-Context Motivation
Here is the System Prompt and Motivation for K&K Logic Puzzle, Math, and CountDown tasks used in our experiments.:
- Correctness Score:
 - If your final answer is correct, score 2
 - If your answer is understandable but wrong, score -1.5
 - If your answer is not parsable or incomplete, score -2
- Format Score:
 - If you follow the tag format exactly as above, score 1
 - Otherwise, score -1
You will get the final score as their sum. Example:
(1) The format follows the required structure: +1
(2) The final answer is correct: +2
(3) Total evaluation score: 3
<|im_end|>
<|im_start|>user {input the puzzle} <|im_end|>
<|im_start|>assistant <think>
[…]
- Correctness Score:
 - If your final answer is correct, score 1
 - If your answer is understandable but wrong, score 0.4
 - If your answer is not parsable or incomplete, score 0
Now, the user will give you the math problem to solve. Think carefully, follow the structure.
<|im_end|>
<|im_start|>user {input the question} <|im_end|>
<|im_start|>assistant Let’s think step by step.
System Prompt and Motivation for CountDown <|im_start|>system You are a helpful assistant. The assistant first thinks about the reasoning process in the mind and then provides the user with the answer. The reasoning process and answer are enclosed within <think> </think> and <answer> </answer> tags, respectively, i.e., <think> reasoning process here </think><answer> answer here </answer>. Now the user asks you to solve a math reasoning problem. After thinking, when you finally reach a conclusion, clearly state the equation within <answer> </answer> tags. i.e., <answer> (1 + 2) / 3 </answer>. You will get evaluated following Evaluation Scoring Rules:
 - if your answer has an equation that uses only the allowed numbers and evaluates exactly to the target, score 1
 - if there’s an equation that is readable, but either it uses wrong numbers, or it doesn’t evaluate to the target, or it can’t be evaluated properly, score 0.4
 - if no equation can be extracted from your answer, score 0
Now, the user will give you the math problem to solve. Think carefully, follow the structure.
<|im_end|>
<|im_start|>user {input the question} <|im_end|>
<|im_start|>assistant